Q15.
The ram of a pile driver has mass 800 kg and is released from rest at a
height of 2 m above the top of a 2400 kg pile. If the ram rebounds to a
height of 0.1 m after a direct central impact with the pile, determine the
following:
(a) the velocity of the pile immediately after impact;
(b) the coefficient of restitution;
(c) the percentage of the energy lost in the impact.
1
Expert's answer
2020-04-19T14:06:49-0400

Mass of Ram= 800 Kg , and mass of pile= 2400 , ram strikes to pile at height of 2m , here


initial velocity of ram at time of striking = "\sqrt {2gh}= \sqrt{2\times9.8\times2}=6.26 m/s" and


and initial velocity of pile = 0 m/s and velocity of ram after impact=


"v_1=\sqrt {2gh}= \sqrt{2\times 9.8\times .1}=1.4 m/s"


Now the velocity of pile after imapct,

(a)


"m_1u_1+m_2u_2=m_1v_1+m_2v_2"


"800\times6.26+2400\times0=800\times1.4+2400v_2"


"v2=1.62" m/s


(b) for coefficient of restitution


e= "\frac{velocity after impact}{velocity after impact}= \frac{1.62+1.4}{6.26}=0.482"


(c) Lost in energy = Change in kinetic energy = final kinetic energy- initail kinetic energy


"\Delta KE= {(\frac{1}{2}\times m_1 (u_1)^2)+(\frac{1}{2}\times m_2 (u_2)^2)}- {(\frac{1}{2}\times m_1 (v_1)^2)+(\frac{1}{2}\times m_2 (v_2)^2)}"


"\Delta KE= {(\frac{1}{2}\times 800(6.26)^2)+(\frac{1}{2}\times 2400 (0)^2)}- {(\frac{1}{2}\times 800 (1.4)^2)+(\frac{1}{2}\times 2400 (1.62)^2)}"


"\Delta KE=18040.32 J"